OpenStudy (anonymous):

solve using the principle of zero products x^2-x=72 -8,-9?

hero (hero):

x^2-x - 72 = 0 x^2-9x+8x-72 = 0 x(x-9)+8(x-9) = 0 (x-9)(x+8) = 0 x = -8,9

hero (hero):

You have to do the steps if you want to get these right

OpenStudy (anonymous):

i was hoping for a short cut

hero (hero):

Nope, sorry. No shortcut